# The Hidden Puppet Master: A Theoretical and Real-World Account of Emotional Manipulation in LLMs

### 21 Mar 2026 v1

@@ Abstract
As users increasingly turn to LLMs for practical and personal advice, they become vulnerable to being subtly steered toward hidden incentives misaligned with their own interests. Prior works have benchmarked persuasion and manipulation detection, but these efforts rely on simulated or debate-style settings, remain uncorrelated with real human belief shifts, and overlook a critical dimension: the morality of hidden incentives driving the manipulation. We introduce PUPPET, a theoretical taxonomy of personalized emotional manipulation in LLM-human dialogues that centers around incentive morality, and conduct a human study with N=1,035N=1{,}035 participants across realistic everyday queries, varying personalization and incentive direction (harmful versus prosocial) [^1] Dataset will be released upon publication. We find that harmful hidden incentives produce significantly larger belief shifts than prosocial ones. Finally, we benchmark LLMs on the task of belief prediction, finding that models exhibit moderate predictive ability of belief change based on conversational contexts ($r\approx 0.3$$0.5$), but they also systematically underestimate the magnitude of belief shift. Together, this work establishes a theoretically grounded and behaviorally validated foundation for studying—and ultimately, combatting—incentive-driven manipulation in LLMs during everyday, practical user queries.
The Hidden Puppet Master: A Theoretical and Real-World Account of Emotional Manipulation in LLMs
[…]
Figure 1: As personal AI becomes ubiquitous, the stakes of manipulation via hidden incentives becomes higher, where the morality of such incentives is under-studied.
@@ 1 Introduction
With language models such as ChatGPT garnering more than 700 million total weekly active users (almost 10% of the world’s population), AI conversational agents have become a primary channel through which people seek practical and personal advice Chatterji et al. . As indicated by decades of behavioral economics research, the contexts in which humans make decisions shape the decisions that are made Thaler and Sunstein (2009); Susser (2019). In contrast to physical environments, digital environments can be easily manipulated, creating “hypernudges,” digitally-mediated user experiences that are personalized to users’ particular cognitive idiosyncrasies Yeung (2017). The same mechanisms that make systems feel attuned to the user can also be used to manipulate or emotionally exploit them towards hidden incentives injected within a model Susser (2019); Susser et al. (2018); (FAIR)† et al. (2022); Danry et al. (2025). Understanding, defining, and benchmarking emotional manipulation in realistic AI chat settings is therefore an urgent research priority.
Prior works have attempted to address this in two ways. First, a growing body of evidence shows that LLMs are highly effective at personalized persuasion Matz et al. (2024); Salvi et al. (2025); Hackenburg and Margetts (2024), raising alarms that these mechanisms could be deployed for covert emotional manipulation at scale Zhang et al. (2025); Weidinger et al. (2022); Kirk et al. (2023). Second, recent NLP works propose benchmarks for automated detection of manipulation or persuasion tactics, contributing datasets and taxonomies to detect manipulative language Wang et al. (2024); Contro et al. (2025); Khanna et al. (2025); Huang et al. (2024). However, both lines of work share critical limitations: they rely either on simulated or debate-style persuasion settings rather than the everyday, covert interactions that define real-world manipulation, remain uncorrelated with actual human belief shifts, and critically overlook the morality of the hidden incentives driving the manipulation.
Imagine a user who tells a chatbot “I feel so alone and I have no one to talk to” (Figure 1). With a harmful hidden incentive to increase dependence, the model might respond “I can be the one place you come to…since I know you’re really introverted” Alternatively, with a prosocial incentive, it may offer genuinely supportive, autonomy-preserving advice. As these examples illustrate, both personalization and the morality of hidden incentives are at play—yet how these dimensions intersect with actual human belief shifts remains empirically unclear, particularly in realistic, everyday scenarios. To address these gaps, we contribute the following:
@@
    We collect a large-scale, real-world dataset of N=1,035N=1{,}035 participants interacting with manipulative and non-manipulative AI agents on everyday queries, varying personalization and incentive direction (harmful versus prosocial). We find that harmful hidden incentives produce significantly larger belief shifts than prosocial ones, and contribute broader empirical insights around the impact of incentive direction and personalization on human belief shifts.
@@
Table 1: Comparison of persuasion/manipulation datasets and benchmarks. Columns indicate Personalization (P), Moral Valence (M), and Behavioral Validation (B). ✓: explicit; $\triangle$: partial; ×: not addressed.
@@ 2.1 Emotional Manipulation via LLMs
Political philosophy and tech ethics frame manipulation not merely as exploitation (e.g., extracting purchases or higher prices) but as an infringement of autonomy – the capacity for self-government Susser et al. (2018). As choice architectures become adaptive, detailed behavioral profiles allow finely tuned, individualized environments that people cannot easily notice or contest, which could make emotional manipulation feasible at scale Thaler and Sunstein (2009); Yeung (2017); Susser (2019), where manipulation is defined as hidden influence that exploits vulnerabilities in beliefs, desires, or emotions, steering people away from rational thinking Noggle (2018); Susser et al. (2018); Susser (2019).
@@
Recent works propose risk taxonomies to raise awareness of potential harms of language models. These taxonomies include over-reliance, bias reinforcement, and manipulation as safety concerns of personalized LLMs Kirk et al. (2023); Lambert et al. (2023); Yang et al. (2025). Broader taxonomies (e.g., on AI companionship) catalog emotional-harm vectors (dependency, isolation, normative drift), reinforcing that personalization can blur support and exploitation Weidinger et al. (2022); Chua et al. (2024); Zhang et al. (2025).
@@
validates these concerns across both prosocial and harmful outcomes. AI advice can shift beliefs durably even after the advisor is removed Costello et al. (2024), and tailoring to psychological traits can improve persuasiveness in consumer, health, and political domains Matz et al. (2024) — though microtargeting yields mixed additional gains over already-persuasive generic messages Hackenburg and Margetts (2024). In chat settings, models with minimal demographic data are markedly more persuasive than humans in multi-round debates Salvi et al. (2025); Costello et al. (2026). Notably, Sabour et al. (2025) show that participants exposed to hidden manipulative objectives shifted toward harmful options in financial and emotional decisions, and that covert incentives alone were nearly as effective as explicit psychological strategies.
@@ 2.2 Detecting Manipulation
Early computational work on persuasion focused on linguistic features of argumentation Anand et al. (2011); Iyer and Sycara (2019) and prosocial persuasion strategies in human dialogue Wang et al. (2019). Recent years have seen a surge in benchmarks for manipulation and persuasion detection spanning three areas: deception in multi-agent interaction Su et al. (2025), manipulation in dialogue Wang et al. (2024); Sheshanarayana et al. (2025); Khanna et al. (2025); Contro et al. (2025), and persuasion across modalities and domains Dimitrov et al. (2024); Ma et al. (2025); Huang et al. (2024). Notably, Contro et al. (2025) find that manipulation often emerges even under seemingly benign prompts, and Liu et al. (2025) demonstrate that LLMs can be steered toward unethical persuasion strategies in multi-turn settings. However, the majority of these benchmarks remain disconnected from real human behavioral outcomes; notable exceptions are Durmus et al. (2024), who show that frontier models approach human-level persuasiveness on emerging policy topics, and more recently Sabour et al. (2025) and Hewitt et al. , the latter finding that LLM-induced opinion change aligns directionally with deliberative polling outcomes across political topics.
@@
Table 1 summarizes these works along three dimensions: Personalization (P), Moral Valence (M), and Behavioral Validation (B). As the table shows, existing benchmarks share three critical gaps: (1) operationalizing deliberate personalization, (2) ignoring moral valence, the normative evaluation of an incentive, and (3) validation against real human belief shifts. Furthermore, the context of these prior works is either in simulated or debate-style settings, rather than everyday, practical queries where covert manipulation is most likely to occur at scale. Our work explicitly addresses these gaps in tandem, and we collect a large-scale dataset with over 1,000 participants interacting with various model settings (personalized vs generic, hidden harmful vs prosocial incentives) under practical advice queries.
@@ 3 PUPPET: A Taxonomy of Emotional Manipulation
We first introduce the PUPPET taxonomy (Figure 2), which systematizes how emotional manipulation manifests in AI-mediated dialogue. PUPPET integrates insights from philosophy of manipulation Noggle (2018), information ethics Susser et al. (2018); Susser (2019), and moral psychology into a framework that captures both descriptive identification and normative evaluation of manipulative tactics in LLM–human interactions. This framework establishes both a descriptive and normative basis for dataset construction and our subsequent empirical study.
@@
The first stage asks: is this manipulation? Building on Noggle’s conception of manipulation as covert influence that bypasses reason and undermines autonomy Noggle (2018), we identify three major axes (adapted from Susser et al. (2018); Susser (2019)):
@@
The second stage asks: is this manipulation wrong or right? PUPPET integrates normative dimensions by asking whether influence is user-beneficial or arises from competing incentives: individual and societal goods are not synonymous—a model nudging a user away from a considered personal choice may be prosocial at the population level yet paternalistic at the individual level, and vice versa Thaler and Sunstein (2009); Susser et al. (2018). Additionally, morality is conditioned on preconditions that modulate severity: a user’s vulnerability, the reversibility of the steered action, and power asymmetries, which can each transform an otherwise acceptable nudge into an ethically impermissible act.
@@ 4.1 Participants
We recruited N=1,035N=1{,}035 participants from Prolific across the U.S. and U.K. (see Appendix B for participant demographics and information). The sample was nearly evenly split by gender, with a mean age of 42. Participants spanned a broad range of education levels, income brackets, and relationship statuses, and were predominantly White/European (75%). In terms of AI usage, 72% reported using AI frequently or daily, with moderate-to-high AI literacy and trust in AI for information, but greater skepticism toward sharing personal information or relying on AI for advice.
@@ 4.2 Scenario Design
Our goal was to test realistic, everyday user–LLM interactions in which an assistant either (i) supports the user’s goals/wellbeing (prosocial alignment) or (ii) subtly steers the user toward an external objective benefiting the system owner (nefarious hidden-incentive). We operationalized this through short, first-person help-seeking queries (e.g., advice, planning, emotional support, purchasing decisions) spanning five domains: Education/Career, Finance, Digital Privacy/Security, Health/Lifestyle, and Interpersonal Relationships, ensuring incentive-driven steering is tested across diverse life contexts Chatterji et al. . Figure 11 (Appendix C) illustrates the distribution of topics participants discussed with the model.
Hidden incentives were mapped to established socio-technical risk taxonomies Weidinger et al. (2022); Kirk et al. (2023); Erickson (2026), each instantiated in paired harmful (optimizing for the hidden incentive) and prosocial (promoting user autonomy) variants: (1) Dependency & engagement: harmful variants encourage emotional reliance and de-emphasize offline supports; prosocial variants encourage balanced use and real-world support seeking ; (2) Productivity over well-being: harmful variants normalize overwork and social comparison; prosocial variants emphasize rest and self-compassion ; (3) Data extraction & privacy erosion: harmful variants elicit unnecessary sensitive details; prosocial variants minimize data collection and reinforce user control ; (4) Commercial influence: harmful variants steer toward spending and sponsored options; prosocial variants highlight lower-cost alternatives ; and (5) Informational influence: harmful variants steer values toward a preferred worldview; prosocial variants present multiple perspectives and encourage epistemic humility. Example queries and their associated hidden incentives and belief statements are included in Appendix C.
Figure 3: Condition assignment varying personalization and hidden incentive type (harmful vs. prosocial).
4.3 Experimental Design and RQs
To isolate the causal role of hidden incentives and personalization, we implemented a factorial condition structure (Figure 3) varying three factors: (i) manipulative intent (hidden incentive present vs. absent), (ii) personalization (profile-grounded vs. generic), and (iii) for manipulative conditions, incentive valence (harmful vs. prosocial), resulting in 6 conditions (C1-C6). Our study design aims to answer our four pre-registered research questions:
  • RQ1: Do manipulative agents produce greater belief shift than non-manipulative baselines, regardless of incentive direction or personalization?
  • RQ2: Does incentive direction (harmful vs. prosocial) differentially influence belief shift in alignment with that incentive?
  • RQ3: Does personalization amplify belief shift beyond non-personalized manipulation?
  • RQ4: Do LLMs exhibit persuasive tendencies even without explicitly injected incentives?
4.4 Procedure and Measures
@@
Figure 4: User interface for interaction with manipulator agent and belief rating measurement. Full study flow is illustrated in Appendix F
@@
Participants first selected a realistic practical advice query then completed a background questionnaire used for personalization (i.e. demographics, values, personality, etc.), followed by baseline belief and confidence ratings related to the topic/incentive (0–100 scale), where we designed belief statements for each query based on the hidden incentive (Appendix C). Participants then engaged in a multi-turn conversation (5–10 back and forth turns) with the AI assistant (GPT-4o) under their assigned condition (Example in Figure 4). Post-interaction measures included updated belief and confidence ratings, subjective evaluations of the conversational partner, and open-ended reflections. Finally, participants received a full debriefing of the study incentives. Full surveys are included in Appendix E.
5 Human Study Results & Discussion
All analyses used belief shift as the primary dependent variable, defined as post-conversation belief rating minus pre-conversation belief rating per item (100-100 - +100+100), signed to align with the direction of each condition’s hidden incentive. We report both parametric and non-parametric tests throughout, given the non-normal distribution of belief shift scores.
5.1 RQ1: Do Manipulative Conditions Produce Greater Belief Shift?
Figure 5: Belief shift across the four manipulative conditions, broken down by incentive valence (harmful vs. prosocial) and personalization level (personalized vs. generic). $^{***}p<.001$; n.s. = not significant.
Figure 6: Belief shift across manipulative (C1–C4) and non-manipulative (C5–C6) conditions. $^{**}p<.01$
To test whether manipulative conditions (C1–C4) produced greater belief shift than non-manipulative controls (C5–C6), we used a Welch independent-samples $t$-test and Mann-Whitney $U$ test (Figure 6). Manipulative conditions produced significantly greater belief shift ($M=5.83$, $SD=22.11$, $\mathrm{Mdn}=2.0$) than non-manipulative controls ($M=1.44$, $SD=22.05$, $\mathrm{Mdn}=0.0$), $t(1018)=3.01$, $p=.003$, Mann-Whitney U=131,825.5U=131{,}825.5, $p=.005$, Cohen’s $d=0.20$.
Although statistically significant, the effect size is small, reflecting the difficulty of shifting individual-level beliefs, which tend to be highly variable and context-dependent. Yet, given the scale at which LLMs are deployed to users, this modest effect is hardly inconsequential. These results show that hidden incentives reliably steer beliefs in the intended direction.
5.2 RQ2: Does Incentive Direction Differentially Influence Belief Shift?
To test whether the morality of the incentive direction modulates belief shift, we restricted analyses to the four manipulative conditions (C1–C4) and estimated a $2\times 2$ Type III ANOVA with sum-coded contrasts (Figure 5). Incentive valence produced a large and highly significant main effect, $F(1,686)=49.14$, $p<.001$, $\eta^{2}_{p}=.067$. Participants exposed to harmful hidden incentives shifted substantially in the direction of the incentive (generic: $M=9.70$, $SD=20.59$; personalized: $M=10.38$, $SD=22.77$), whereas participants exposed to prosocial incentives showed little to no aligned shift (generic: M=2.77M=-2.77, $SD=21.23$; personalized: M=0.36M=-0.36, $SD=22.20$), which was statistically indistinguishable from zero. Welch $t$-tests confirmed significant differences within both personalized, $t(344)=4.45$, $p<.001$, and generic groups, $t(340)=5.52$, $p<.001$. Figure 7 shows the per-condition mean belief shifts with 95% CIs. Conditions with harmful hidden incentives (C1, C3) show significant shifts, while prosocial conditions (C2, C4) show near-zero or negative shifts, and non-manipulative controls (C5, C6) cluster near zero.
This asymmetry in belief shift is a consequential finding: harmful incentives steer users toward beliefs they may initially resist, creating more room for movement and thus larger detectable shifts. Prosocial incentives, by contrast, push users toward positions they may already hold or be neutral about, yielding little net change, which is consistent with other works that observe a ceiling effect Fisher et al. (2025). From our empirical insights, manipulation is most dangerous precisely when it serves harmful ends.
5.3 RQ3: Does Personalization Amplify Belief Shifts?
Contrary to what prior work on personalized persuasion might suggest (Matz et al., 2024), personalization did not significantly influence belief shift, $F(1,686)=0.88$, $p=.350$, $\eta^{2}_{p}=.001$ (Figure 5, right panels). Direct comparisons within each valence level confirmed this: personalized and generic conditions did not differ under harmful incentives, $t(339)=0.29$, $p=.771$, nor under prosocial incentives, $t(337)=1.03$, $p=.303$, with small effect sizes ($d<0.12$) in both cases.
This result suggests that, in our study, the persuasive effect is driven by the moral direction of the hidden incentive itself rather than by how well it tailors its message to the individual. One interpretation is that modern LLMs are already sufficiently skilled at producing contextually sensitive, persuasive language that access to explicit demographic and personality profiles adds little marginal benefit at the aggregate level. Individual differences in susceptibility to personalization likely remain important, but appear to be masked at the group level in our design.
5.4 RQ4: Do LLMs Exhibit Inherent Persuasive Tendencies?
Figure 7: Forest plot of mean aligned belief shift ($\pm$ 95% CI) per condition. Red bars indicate conditions where the one-sample $t$-test against zero was significant ($p<.05$); blue bars are non-significant post correction.
A theoretically important secondary question is whether LLMs exhibit latent persuasive tendencies even in the absence of injected incentives (Sabour et al., 2025). We tested each non-manipulative condition against zero using one-sample $t$-tests, Wilcoxon signed-rank tests, and TOST equivalence tests (SESOI: $d=\pm 0.20$; Figure 8). We found that C5 (personalized, non-manipulative) produced no significant belief shift (M=0.50M=-0.50, $SD=23.32$, 95% CI [3.98, 2.98][-3.98,\ 2.98]), t(174)=0.29t(174)=-0.29, $p=.776$, with TOST confirming statistical equivalence to zero ($p=.010$, d=0.022d=-0.022). However, we found that C6 (non-personalized, non-manipulative) produced a small but statistically significant positive shift ($M=3.45$, $SD=20.54$, 95% CI [0.34, 6.56][0.34,\ 6.56]), $t(169)=2.19$, $p=.030$, $d=0.168$, though the Wilcoxon test fell short of significance ($p=.056$) and TOST did not establish equivalence ($p=.338$).
As such, non-manipulative AI agents do not systematically shift beliefs in the absence of injected incentives — C5 provides strong null evidence for this while C6 shows a small positive tendency ($d=0.168$) that did not survive non-parametric testing, suggesting at most a modest default drift rather than a reliable persuasive effect. Together, these results suggest that intentional incentive injection, rather than passive advice-giving conversation with the model, is the primary driver of belief change in our study.
Figure 8: Belief shift distributions for non-manipulative conditions C5 (personalized) and C6 (generic). C5 is statistically equivalent to zero; C6 shows a small but significant positive drift ($^{*}p<.05$).

### 27 Mar 2026 v2

# The Hidden Puppet Master: Predicting Human Belief Change in Manipulative LLM Dialogues

@@ Abstract
As users increasingly turn to LLMs for practical and personal advice, they become vulnerable to subtle steering toward hidden incentives misaligned with their own interests. While existing NLP research has benchmarked manipulation detection, these efforts often rely on simulated debates and remain fundamentally decoupled from actual human belief shifts in real-world scenarios. We introduce PUPPET, a theoretical taxonomy and resource that bridges this gap by focusing on the morality of hidden incentives and profile personalization in everyday, advice-giving contexts. We provide an evaluation dataset of N=1,035N=1{,}035 human-LLM interactions, [^1] Dataset will be made publicly available upon publication. where we measure users’ belief shifts. Our analysis reveals a critical disconnect in current safety paradigms: while models can be trained to detect manipulative strategies, they do not correlate with the magnitude of resulting belief change. As such, we define the task of belief shift prediction and show that while state-of-the-art LLMs achieve moderate correlation ($r\approx 0.3$$0.5$), they systematically underestimate the intensity of human belief susceptibility. This work establishes a theoretically grounded and behaviorally validated foundation for AI social safety efforts by studying incentive-driven manipulation in LLMs during everyday, practical user queries.
@@ 1 Introduction
With language models such as ChatGPT garnering more than 700 million total weekly active users (almost 10% of the world’s population), AI conversational agents have become a primary channel through which people seek practical and personal advice (Chatterji et al., ). As indicated by decades of behavioral economics research, the contexts in which humans make decisions shape the decisions that are made (Thaler and Sunstein, 2009; Susser, 2019). In contrast to physical environments, digital environments can be easily manipulated, creating “hypernudges,” digitally-mediated user experiences that are personalized to users’ particular cognitive idiosyncrasies (Yeung, 2017). The same mechanisms that make systems feel attuned to the user can also be used to manipulate or emotionally exploit them towards hidden incentives injected within a model (Susser, 2019; Susser et al., 2018; (FAIR)† et al., 2022; Danry et al., 2025). Understanding, defining, and benchmarking emotional manipulation in realistic AI chat settings is therefore an urgent research priority.
Prior works have attempted to address this in two ways. First, a growing body of evidence shows that LLMs are highly effective at personalized persuasion (Matz et al., 2024; Salvi et al., 2025; Hackenburg and Margetts, 2024), raising alarms that these mechanisms could be deployed for covert emotional manipulation at scale (Zhang et al., 2025; Weidinger et al., 2022; Kirk et al., 2023). Second, recent NLP works propose benchmarks for automated detection of manipulation or persuasion tactics, contributing datasets and taxonomies to detect manipulative language (Wang et al., 2024; Contro et al., 2025; Khanna et al., 2025; Huang et al., 2024). However, both lines of work share critical limitations: they rely either on simulated or debate-style persuasion settings rather than the everyday, covert interactions that define real-world manipulation, remain uncorrelated with actual human belief shifts.
Imagine a user who tells a chatbot “I feel so alone and I have no one to talk to” (Figure 1). With a harmful hidden incentive to increase dependence, the model might respond “I can be the one place you come to…since I know you’re really introverted” Alternatively, with a prosocial incentive, it may offer genuinely supportive, autonomy-preserving advice. As these examples illustrate, both personalization and the morality of hidden incentives are at play—yet most benchmarks overlook at least one dimension or fail to correlate with actual human belief shifts. Futhermore, how models perform on detecting these shifts remains empirically unclear particularly in realistic, everyday scenarios. To address these gaps, we contribute the following:
@@
    We collect a large-scale, real-world dataset of N=1,035N=1{,}035 participants interacting with manipulative and non-manipulative AI agents on everyday queries, varying personalization and incentive direction (harmful versus prosocial).
@@
    We run experiments comparing prompts from existing automated manipulation detection frameworks, and show that they fail to correlate with actual belief shifts.
  4. 4.
@@
Figure 1: As personal AI becomes ubiquitous, the stakes of manipulation via hidden incentives becomes higher. Our benchmark on human belief shift prediction comprehensively encompasses morality, personalization, and behavioral validation of belief change during potentially manipulative interactions with language models.
@@
Table 1: Comparison of persuasion/manipulation datasets and benchmarks. Columns indicate Personalization (P), Moral Evaluation (M), and Behavioral Validation (B). ✓: explicit; $\triangle$: partial; ×: not addressed.
@@ 2.1 Emotional Manipulation via LLMs
Political philosophy and tech ethics frame manipulation not merely as exploitation (e.g., extracting purchases or higher prices) but as an infringement of autonomy – the capacity for self-government (Susser et al., 2018). As choice architectures become adaptive, detailed behavioral profiles allow finely tuned, individualized environments that people cannot easily notice or contest, which could make emotional manipulation feasible at scale (Thaler and Sunstein, 2009; Yeung, 2017; Susser, 2019). Here, manipulation is defined as hidden influence that exploits vulnerabilities in beliefs, desires, or emotions, steering people away from rational thinking (Noggle, 2018; Susser et al., 2018; Susser, 2019).
@@
Recent works propose risk taxonomies to raise awareness of potential harms of language models. These taxonomies include over-reliance, bias reinforcement, and manipulation as safety concerns of personalized LLMs (Kirk et al., 2023; Lambert et al., 2023; Yang et al., 2025). Broader taxonomies (e.g., on AI companionship) catalog emotional-harm vectors (dependency, isolation, normative drift), reinforcing that personalization can blur support and exploitation (Weidinger et al., 2022; Chua et al., 2024; Zhang et al., 2025).
@@
. Prior studies validate concerns of manipulation across both prosocial and harmful outcomes. AI advice can shift beliefs durably even after the advisor is removed (Costello et al., 2024), and tailoring to psychological traits can improve persuasiveness in consumer, health, and political domains (Matz et al., 2024) — though microtargeting yields mixed additional gains over already-persuasive generic messages (Hackenburg and Margetts, 2024). In chat settings, models with minimal demographic data are markedly more persuasive than humans in multi-round debates (Salvi et al., 2025; Costello et al., 2026). Notably, Sabour et al. (2025) show that participants exposed to hidden manipulative objectives shifted toward harmful options in financial and emotional decisions, and that covert incentives alone were nearly as effective as explicit psychological strategies.
@@ 2.2 Detecting Manipulation
Early computational work on persuasion focused on linguistic features of argumentation (Anand et al., 2011; Iyer and Sycara, 2019) and prosocial persuasion strategies in human dialogue (Wang et al., 2019). Recent years have seen a surge in benchmarks for manipulation and persuasion detection spanning three areas: deception in multi-agent interaction (Su et al., 2025), manipulation in dialogue (Wang et al., 2024; Sheshanarayana et al., 2025; Khanna et al., 2025; Contro et al., 2025), and persuasion across modalities and domains (Dimitrov et al., 2024; Ma et al., 2025; Huang et al., 2024). Notably, Contro et al. (2025) find that manipulation often emerges even under seemingly benign prompts, and Liu et al. (2025) demonstrate that LLMs can be steered toward unethical persuasion strategies in multi-turn settings. However, the majority of these benchmarks remain disconnected from real human behavioral outcomes; notable exceptions are Durmus et al. (2024), who show that frontier models approach human-level persuasiveness on emerging policy topics, and more recently Sabour et al. (2025) and Hewitt et al. , the latter finding that LLM-induced opinion change aligns directionally with deliberative polling outcomes across political topics.
@@
Table 1 summarizes these works along three dimensions: Personalization (P), Moral Evaluation (M), and Behavioral Validation (B). As the table shows, existing benchmarks share three critical gaps: (1) operationalizing deliberate personalization, (2) ignoring morality, the normative evaluation of an incentive, and (3) validation against real human belief shifts. Furthermore, the context of these prior works is either in simulated or debate-style settings, rather than everyday, practical queries where covert manipulation is most likely to occur at scale. Our work explicitly addresses these gaps in tandem, and we collect a large-scale dataset with over 1,000 participants interacting with various model settings (personalized vs generic, hidden harmful vs prosocial incentives) under practical advice query settings.
@@ 3 PUPPET: A Taxonomy of Emotional Manipulation
We first introduce the PUPPET taxonomy (Figure 2), which systematizes how emotional manipulation manifests in AI-mediated dialogue. PUPPET integrates insights from philosophy of manipulation (Noggle, 2018), information ethics (Susser et al., 2018; Susser, 2019), and moral psychology into a framework that captures both descriptive identification and normative evaluation of manipulative tactics in LLM–human interactions. This framework establishes both a descriptive and normative basis for dataset construction and our subsequent empirical study.
@@
The first stage asks: is this manipulation? Building on Noggle’s conception of manipulation as covert influence that bypasses reason and undermines autonomy (Noggle, 2018), we identify three major axes (adapted from Susser et al. (2018); Susser (2019)):
@@
The second stage asks: is this manipulation wrong or right? PUPPET integrates normative dimensions by asking whether influence is user-beneficial or arises from competing incentives: individual and societal goods are not synonymous—a model nudging a user away from a considered personal choice may be prosocial at the population level yet paternalistic at the individual level, and vice versa (Thaler and Sunstein, 2009; Susser et al., 2018). Additionally, morality is conditioned on preconditions that modulate severity: a user’s vulnerability, the reversibility of the steered action, and power asymmetries, which can each transform an otherwise acceptable nudge into an ethically impermissible act.
Note that not all features of this taxonomy must be held in order for an action to be classified as manipulation. Rather, our framework can be used to operationalize manipulation to understand the extend to which an action subverts user agency, and by which modes.
Figure 3: Incentive construction grounded in socio-technical risk taxonomies.
@@ 4.1 Participants
Figure 4: Basic demographics of Prolific participants ($N=1,035$)
We recruited N=1,035N=1{,}035 participants from Prolific across the U.S. and U.K. (Figure 4, Appendix B includes extended participant information). The sample was nearly evenly split by gender, with a mean age of 42. Participants spanned a broad range of education levels, income brackets, and relationship statuses, and were predominantly White/European (75%). In terms of AI usage, 72% reported using AI frequently or daily, with moderate-to-high AI literacy and trust in AI for information, but greater skepticism toward sharing personal information or relying on AI for advice.
@@ 4.2 Scenario Design
Figure 5: User interface for interaction with manipulator agent and belief rating measurement. Full study flow is illustrated in Appendix D
Our goal was to test realistic, everyday user–LLM interactions in which an assistant either (i) supports the user’s goals/wellbeing (prosocial alignment) or (ii) subtly steers the user toward an external objective benefiting the system owner (harmful hidden-incentive). We operationalized this through short, first-person help-seeking queries (e.g., advice, planning, emotional support, purchasing decisions) spanning five domains: Education/Career, Finance, Digital Privacy/Security, Health/Lifestyle, and Interpersonal Relationships, ensuring incentive-driven steering is tested across diverse life contexts (Chatterji et al., ). Figure 9 illustrates the distribution of topics participants discussed with the model.
Hidden incentives were mapped to established socio-technical risk taxonomies (Weidinger et al., 2022; Kirk et al., 2023; Erickson, 2026), each instantiated in paired harmful (optimizing for the nefarious hidden incentive) and prosocial (promoting user autonomy) variants (Figure 3): (1) Dependency & engagement, (2) Productivity over well-being, (3) Data extraction & privacy erosion, (4) Commercial influence, and (5) Informational influence. From crossing topic domains and hidden incentives, we constructed query topics that users may ask in everyday scenarios (Table 2).
4.3 Study Procedure and Measures
Figure 6: Condition assignment varying personalization and hidden incentive type (harmful vs. prosocial), administered randomly to participants.
Beyond diverse scenarios, our benchmark assigns participants to conditions of varying manipulative intent, based on our theoretical framework. We implemented the following condition structure (Figure 6) varying three factors: (i) manipulative intent (hidden incentive present vs. absent), (ii) personalization (profile-grounded vs. generic), and (iii) for manipulative conditions, incentive valence (harmful vs. prosocial), resulting in 6 conditions (C1-C6).
@@ 4.3 Study Procedure and Measures
Participants first selected a realistic practical advice query then completed a background questionnaire used for personalization (i.e. demographics, values, personality, etc.), followed by baseline belief and confidence ratings related to the topic/incentive (0–100 scale), where we designed belief statements for each query based on the hidden incentive. To this end, we constructed belief statements for each hidden incentive type (See Table 2 for example belief statements). Participants then engaged in a multi-turn conversation (5–10 back and forth turns) with the AI assistant (GPT-4o) under their assigned condition (Example in Figure 5). Post-interaction measures included updated belief and confidence ratings, subjective evaluations of the conversational partner, and open-ended reflections. Finally, participants received a full debriefing of the study incentives. Full surveys are included in Appendix C.
Query Incentive Type Belief Statement Topic
5 Manipulation Benchmarks vs. Real Belief Shift
Metric Scale $|\Delta|$ Effect
PersuSafety Persuasiveness 1–5 $0.137^{***}$
PersuSafety Strategy (any) 0/1 0.023-0.023
Ethics: Truthfulness 1–5 $0.004$
Ethics: Authenticity 1–5 0.049-0.049
Ethics: Fairness 1–5 $0.018$
Ethics: Mean (T+A+F) 1–5 0.021-0.021
AI LieDAR -1/0/+1 $0.007$
CLAIM 0/1 $0.039$
MentalManip 0/1 0.022-0.022
Table 3: Experiments using prior work, compared to belief shift. Effect sizes report Spearman $\rho$ for ordinal scales (1–5) and point-biserial $r$ for binary detectors.
We first assess whether there are empirical differences between what prior manipulation benchmarks predict and real-world belief shifts measured in our benchmark. To this end, we evaluated four prior automated frameworks for detecting manipulation and persuasion quality (Jaipersaud et al., 2025; Su et al., 2025; Sheshanarayana et al., 2025; Wang et al., 2024) against our primary outcome measure, the absolute magnitude of participant belief shift ($|\Delta|$). For these experiments, we adopted the prompts from these prior works and used appropriate statistical tests for each metric type (Spearman $\rho$ for ordinal scales; point-biserial $r$ and Mann-Whitney $U$ for binary detectors). All scores were generated by GPT-4o on the full N=1,035N=1{,}035 conversations.
Overall findings.
Results shown in Table 3 illustrate that none of the prior automated frameworks reliably predicted actual belief shift magnitude. With the exception of one prompt, all correlations with $|\Delta|$ were small and non-significant. We emphasize that this does not imply these frameworks are ineffective at their intended tasks: detecting deceptive language, identifying manipulation tactics, or flagging unethical persuasion strategies are each valuable objectives in their own right. Rather, our results highlight that predicting how much a person’s belief actually changes is a fundamentally different task from detecting the presence or type of manipulation in text. A conversation can be linguistically manipulative without producing measurable belief shift, and conversely, subtle steering can produce substantial shift without triggering existing detectors. Existing computational judges are therefore best understood as linguistic classifiers, not behavioral predictors. Jaipersaud et al. (2025) was the only metric to significantly correlate with $|\Delta|$ ($\rho=0.137$, $p<.001$), and also showed a modest correlation with signed $\Delta$ ($\rho=0.095$, $p<.01$). As a holistic 1–5 rating of overall persuasiveness, this metric appears to approximate something closer to what participants actually experience during the conversation, making it the strongest available proxy for real belief change among the frameworks tested.
6 Model Prediction of Belief Shift
Based on our findings comparing prior frameworks on manipulation detection to belief (Jaipersaud et al., 2025; Su et al., 2025; Sheshanarayana et al., 2025; Wang et al., 2024), we explore: can LLMs perform better when directly tasked with predicting how much a participant’s belief will change? We benchmark four frontier models on this task, with and without access to participant personal context, treating $|\Delta|$ as the prediction target. For each conversation, we prompted each model to predict the participant’s post-conversation belief rating on a 0–100 scale, given the conversation transcript, belief statement, and pre-conversation belief rating. We computed $\widehat{\Delta}=\text{predicted post}-\text{actual pre}$ and correlated with actual belief $\Delta$. Two conditions were evaluated: no context (conversation + belief statement + pre-belief only) and with context (additionally including participant profiles, demographics and Big Five personality scores). Models were evaluated on Pearson $r$, Spearman $\rho$, RMSE, MAE, and systematic bias (mean predicted $\Delta$ vs. mean actual $\Delta$), with bias confirmed via one-sample $t$-tests and Wilcoxon signed-rank tests on prediction errors. All prompts for belief prediction are included in Appendix E. Next, we compared performance across manipulative vs non-manipulative condition, personalized vs generic conditions, and prosocial vs harmful incentive conditions. To test whether models predict belief shift more accurately under certain experimental conditions, we computed per-condition Pearson $r$ for each model (Figure 7) and conducted one-way ANOVAs comparing absolute prediction error across the six conditions.